OpenStudy (anonymous):

For which equation would you use the operations of subtraction and multiplication to solve? five times u minus five equals five the quantity u divided by five all plus five equals five five times u plus five equals five the quantity u divided by five all minus five equals five @mikurout

OpenStudy (mikurout):

option b

OpenStudy (anonymous):

the last one

OpenStudy (anonymous):

Show you're work.

OpenStudy (anonymous):

Idk which one is right.

OpenStudy (mikurout):

option b . according to this u/5 + 5 = 5 to solve this you have to do 1st substraction so u/5 = 5-5 so u/5 = 0 then multiplication so u = 5*0 u = 0

OpenStudy (anonymous):

@jack1 who's right?

OpenStudy (anonymous):

yes sorry i meant option b aswell

OpenStudy (jack1):

@Freelance1 b is correct